5 years ago, John was twice as old as his brother jim. 3 years from nw the sum of thier ages will be 31. how old is jim now?
Answer by sachi(548)

5 years ago, John was twice as old as his brother jim. 3 years from nw the sum of thier ages will be 31. how old is jim now?
ans:-
let jim is x yrs now
5 yrs ago jim was x-5 yrs & then john was 2(x-5)yrs
3 yrs from now jim will be x+3 & john will be 2(x-5)+8
as per question x+3+2(x-5)+8=31
or 2x+3-10+8=31
or x=31-1=30 ans